y= sqrt. x+2 / sqrt. x+4

Grade:

4 Answers

Aman Bansal
11 years ago

Dear Hessica,

 

Domain is x>-2.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Chetan Mandayam Nayakar
11 years ago

domain is x≥-2 AND x<-4

Jessica de vera
11 years ago

what is the range? and how can i find it? thanks in advance

shikhar srivastav
11 years ago

sqrt(x+4)=0

x not equal -4, x+4>=0 ,x>=-4,

x+2>=0,x>=-2

domain is (-2,infinity)   21:08:36